one positive number is 2/3 of the other and their product is 24. what is the sum of the two?
Answer by TimothyLamb(4379)

---
x = (2/3)y
xy = 24
---
xy = 24
(2/3)yy = 24
(2/3)yy - 24 = 0
(0.666666)yy - 24 = 0
---
the above quadratic equation is in standard form, with a=0.666666, b=0, and c=-24
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
0.666666 0 -24
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
this quadratic has two real roots (the x-intercepts), which are:
y = 6
y = -6
---
the negative root does not fit the problem statement, so use the positive root
---
answer:
y = 6
x = 4
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php